2. Estimate Northern and Southern hemispheric average OH concentrations by setting up a box model utilizing the following information Southern Hemisphere Northerrn Hemisphere CH3CCla CH3CCl3 interhemispheric transport OH OH Industrial emissions Industrial emissions
Answer #1

Answer:

Northen hemisphere (-OH) concentration:

Given

CH3CCl3 → OH

MCH3CCl3 = 133.4 g/mol

Volume of atmosphere = 5.1*1024 cm3

Global CH3CCl3 production = 5.8*1011 g/yr

Northen hemisphere production = 95% of total emissions

95%of the total productions = (95/100)*5.8*1011 gr/yr = 5.51*10 11 g/yr

kOH = 6.7*10-15 cm3 moles-1 s-2

OH = (5.51*1011)/(6.7*10-15)

= 0.822 *1026

Out of which (0.781*1026) will produce in northen hemispehere.